Calculate the potential of the following concentration cell at 25.0 ◦C.

Ag | Ag+(1.00×10−3 M) || Ag+(0.100 M) | Ag

Identify the anode and the cathode in the cell. How would the cell potential change if the concentrations
of Ag+ in both half-cells were doubled? Use the Nernst equation to support your answer.
1
Expert's answer
2020-11-10T14:02:44-0500

Ag | Ag+(1.00×10−3 M) || Ag+(0.100 M) | Ag


concentration cell at 25.0 ◦C.

An electrochemical cell of this type, in which the anode and cathode compartments are identical except for the concentration of a reactant, is called a concentration cell. As the reaction proceeds, the difference between the concentrations of Ag+ in the two compartments will decrease, as will Ecell.

2Ag+(aq) +2e- > 2Ag(s)

Cu(s) > Cu2+(aq) + 2e-


Eo reduction = +0.799v

Eo Oxidation = - 0.518v


2Ag+(aq) + Al(s) > 2Ag(s) + Cu2+(aq) EO =+0.281v

N = 2 moles of electrons


Eocell = (0.0592/n) log K


0.281 = (0.0592/2) log K


Log k = 9.49


K = 109.49 =3.1 x 109